Design a Quarter-wave Transformer using a Microstrip Line to Match a 250 Ohm Load to a 75 Ohm transmission Line. The PCB is made of FR-4 and has a thickness of 0.2 mm. Specify Length and width of the strip for a frequency of 2.4 Ghz.
Answer #1

We know that characteristics equation impedance (20) is given by- 5-98 be 0.8w+t - 2o= 87 en ... Ver #tui l given t-2-4 t=0.2put values in eq -o- 136.93= 87 don 5.98X31025 Ertlul (0.8wtv.2 / for air ine Er=1 so 136.93= 87 en 186.875 - 2.ui Con2toash,